Calculate the molar mass of the solute in a solution of 2.47 g of a mo electrolyte in 100 g of acetic acid (CH3COO). The solution freezes at 1.3 °C below the normal freezing point of pure acetic acid. (Kf for CH3COOH=3.9°C/m)
1
Expert's answer
2020-10-26T13:55:03-0400

∆T=KfCm

Cm=∆T/Kf=1.3/3.9=0.3(mol/kg)

Cm=nsolute/msolvent

nsolute=Cmmsolvent=0.3•0.1=0.03(mol)

M=m/n=2.47/0.03=82.33(g/mol)